OpenStudy (anonymous):

Let f(x) = 3 x^2 - 8. Find an equation of the line tangent to the graph of f at the point where x = -3. y =

OpenStudy (mertsj):

Do you know that the first derivative is the tangent function?

OpenStudy (anonymous):

the power rule?

OpenStudy (mertsj):

Yes

OpenStudy (anonymous):

f'(X)= rx^r-1

OpenStudy (mertsj):

f'(x)=6x

OpenStudy (anonymous):

that the deriv 6x so plug that in the equation

OpenStudy (mertsj):

At x = -3, f'(x)=-18 and that is the slope of the tangent at x = -3

OpenStudy (anonymous):

y=mx+b right? or y-y1=mx-x1

OpenStudy (mertsj):

y=6x That is the equation of the tangent function. At x = -3, it has a value of -18 and that is the slope of the tangent at that point.

OpenStudy (mertsj):

Why don't you draw the graph of the parabola and then draw the tangent at x = -3 and see what it looks like?

OpenStudy (mertsj):

That would help you understand.

OpenStudy (anonymous):

wait what is the full equation y=6x?

OpenStudy (mertsj):

So if you want the equation of the tangent you know its slope is -18 and you know it goes through the point (-3,19) So use either y= mx+b or point slope form. Whichever you prefer.

OpenStudy (mertsj):

y=6x is the function that gives you the slope of the tangent at any x value of the parabola.

OpenStudy (anonymous):

y=-18(-3)+b

OpenStudy (mertsj):

\[y-19=-18(x+3)\] \[y=-18x-35\]

OpenStudy (mertsj):

y=-18(-3)+b 19=-18(-3)+b

OpenStudy (anonymous):

for some reason im not writing it corrrectly bcause my book says its wrong

OpenStudy (mertsj):

what does your book say it is?

OpenStudy (anonymous):

equation of the tangent line of y = x3 at x = 2

OpenStudy (anonymous):

3 x^2 = f(x)+8

OpenStudy (mertsj):

We are not working with the function y = x^3. We are doing y = 3x^2-8

OpenStudy (mertsj):

What does your book say for this one?

OpenStudy (anonymous):

it says y=-18x+b

OpenStudy (mertsj):

For the final answer?

OpenStudy (anonymous):

yeah they only want the equation

OpenStudy (anonymous):

of a line

OpenStudy (mertsj):

Well it would be easy to find b and if i were you I would give both answers.

OpenStudy (anonymous):

ok thanks

OpenStudy (mertsj):

Because the problem specifies to find the equation of the tangent at x = -3.

OpenStudy (mertsj):

yw